# Reinforcement Fine-Tuning of Large Language Models

### v1

@@ 3.3.1 End-to-end Data Transformation
  • The Formatter Module unifies disparate data sources into RFT-compatible formats, providing convenient conversion between raw inputs (e.g., meta-prompts, domain-specific corpora, and QA pairs with tagged rewards) and structured RFT representations (e.g., via RftDatasetModel). For efficient RFT workloads, we utilize PostgreSQL-based persistent storage supporting both ExperienceModel for prioritized trajectories and DPODataModel for preference pairs, which is particularly beneficial for asynchronous RFT mode and distributed environments. Full schema details appear in Appendix B.1.
  • The Controller Module manages the complete data pipeline lifecycle through distributed server initialization, declarative configuration, and automated dataset persistence. It implements dynamic control mechanisms for asynchronous scenarios and protection against resource exhaustion, with configurable termination conditions based on compute quota or data quantity. This modular design enables Trinity-RFT to handle data transformations flexibly while maintaining consistency across different RFT modes. More service utilities are detailed in Appendix B.2.
@@
Figure 2: An illustration of data flow in Trinity-RFT.
3.3.2 Active Data-Reward Co-Shaping
To enable efficient and effective learning from complex and delayed reward signals, Trinity-RFT introduces fruitful mechanisms that actively shape both data and reward throughout the RFT process.
@@ 3.3.2 Active Data-Reward Co-Shaping
  • Trinity-RFT provides the Processor Module to support convenient and extensive data processing capabilities tailored to RFT, through dedicated bridge classes that link Trinity-RFT to over 100 operators from the Data-Juicer system [2]. This module offers efficient integration with the experience buffer, and provides composable building blocks for experience cleaning (e.g., length filters, duplication removal), safety alignment (e.g., toxicity detection, ethics checks), and preference data synthesis (e.g., critique-conditioned augmentation). Users can specify processing chains using either natural language or explicit configurations. These are accessible to both RFT users familiar with Data-Juicer and those who are not, thanks to transparency of the underlying system details. By treating Data-Juicer as a modular data processing operator pool rather than a central dependency, Trinity-RFT provides RL-specific abstractions and coherence, while benefiting from well-established data tools.
  • RL-oriented data shaping. RL is particularly useful in scenarios where optimization goals cannot be clearly defined, leading users to express vague or high-level requirements. In response, Trinity-RFT automates the translation of these objectives into optimized data processing pipelines. It reduces problem complexity through a structured hierarchy: Quality $\rightarrow$ Difficulty $\rightarrow$ Diversity $\rightarrow$ Quantity, with auto-selection of appropriate operators from Data-Juicer. Moreover, Trinity-RFT provides many composable and extensible interfaces for data shaping, aiming to seamlessly inject user-defined inductive biases. For instance, DataCleaner (a sub-module of Processor) supports multi-way cleaning with statistical filtering (e.g., entropy-based diversity [11]) and rule-based validation (e.g., boxed answer verification in math dataset [7]), while DataSynthesizer enables both autonomous generation and hybrid augmentation with controlled feedback metrics [35].
  • Active data optimization. As agents interact with environments, experiences collected and stored in the buffer can become large-scale and highly diverse. Using the right experiences for RL training is crucial for a cost-effective and stable learning process. To this end, Trinity-RFT allows users to flexibly apply multi-dimensional utility scoring to prioritize samples (e.g., 0.4 * difficulty + 0.3 * diversity + 0.1 * quality - 0.2 * frequency). Moreover, we introduce the DataActiveIterator module to enable (1) version-controlled experience reuse through ExperienceModel; (2) cross-task data lineage tracking via rich relationships derived from sample_id and rollout-model_id; and (3) incorporating human annotations into the rewards. These features will be demonstrated through concrete examples in Section 4 and Appendix B.3.
@@ 3.3.3 Human-AI Collaboration
The decoupled design of Trinity-RFT, or the presence of an experience buffer in particular, enables human feedback to participate in RL loops without breaking the asynchronous execution model. For instance, human-verified samples can be prioritized for training while fresh experiences are being collected, which is a critical capability for real-world deployment scenarios with mixed feedback sources. Further details for human-AI collaboration in Trinity-RFT will be illustrated in Section 4.4.
3.4 Performance Optimizations and User-Friendliness
Trinity-RFT has made numerous efforts in pursuit of efficiency, for example:
  • For multi-turn conversations and ReAct-style workflows [37], Trinity-RFT supports concatenating multiple rounds of agent-environment interaction compactly into a single sequence, with proper masking that indicates which tokens need to be incorporated into the training objective of RL algorithms.
  • For the synchronous RFT mode, Trinity-RFT accelerates the RFT process via pipeline parallelism between the explorer and trainer.
  • Trinity-RFT utilizes Ray [14] for distributed runtime, and NCCL communication primitives [16] (whenever feasible) for model weight synchronization between the explorer and trainer.
In addition, Trinity-RFT has been designed with user-friendliness as a top priority, for example:
@@ 3.4 Performance Optimizations and User-Friendliness
  • We include a monitor (built upon Wandb [32] and TensorBoard [29]) that allows the user to conveniently track the progress of an RFT process, both quantitatively (e.g., via learning curves for rewards and other metrics) and qualitatively (e.g., via concrete examples of rollout trajectories generated at different RL steps). See Figure 3 for an example snapshot of the monitor.
  • We also implement a configuration manager, as shown in Figure 4, that allows the user to create configuration files conveniently via a front-end interface.
  • We provide Trinity-Studio, an all-in-one unified UI that allows the user to configure and run data inspection, data processing, RFT learning process, etc., all by clicking the mouse and filling forms, without writing any code. More functions will be integrated in the future. An example for developing with Trinity-Studio will be introduced in Section 4.5.
@@
4 Examples and Applications
Through some concrete examples, we demonstrate how to use Trinity-RFT for different purposes, and also exemplifies some concepts introduced in previous sections. Additional step-by-step tutorials can be found on the documentation website [^1] https://modelscope.github.io/Trinity-RFT, or the /examples folder of the GitHub repository [^2] https://github.com/modelscope/Trinity-RFT.
4.1 Single-Turn Workflow
In a simple yet common scenario, a user of Trinity-RFT would like to train an LLM for completing single-turn tasks, where the LLM generates one response to each input query. For this purpose, the user mainly needs to (1) define and register a single-turn workflow class (by inheriting the base class Workflow) tailored to the targeted tasks, and (2) specify the dataset of tasks (for training and/or evaluation) and the initial LLM, both compatible with HuggingFace [9] and ModelScope [13] formats.
Listing 1 gives a minimal example for implementing a single-turn workflow. Suppose that each task is specified by a <question, answer> tuple. The run() method of ExampleWorkflow calls the LLM once to generate a response for the question, calculates its reward, and returns an Experience instance that consists of the response itself, the reward value, and the log-probabilities of next-token prediction by the rollout model (which is necessary for certain RL algorithms, such as PPO [22] and GRPO [23]). Some built-in workflows have been implemented in Trinity-RFT, e.g., the MathWorkflow class for math-related tasks.
Trinity-RFT also allows more flexible and advanced usage by changing just a few configuration parameters, e.g., running supervised fine-tuning (SFT) as a warmup stage before the RFT process, or choosing between the synchronous and asynchronous modes of RFT (as described in Section 3.1).
@@ 4.1 Single-Turn Workflow
4 def __init__(self, model: ModelWrapper, task: Task, **kwargs):
5 super().__init__(model, task, **kwargs)
@@
8
9 def calculate_reward(self, response: str, truth: str) -> float:
@@
11
@@
20 n=self.task.rollout_args.repeat_times,
21 temperature=self.task.rollout_args.temperature,
@@
23 reward: float = self.calculate_reward(response.response_text, self.answer)
@@ 25 Experience(
4.2 Multi-Turn Workflow
In more advanced cases, the user would like to train an LLM-powered agent that solves multi-turn tasks by repeatedly interacting with the environment. With Trinity-RFT, achieving this is mostly as simple as in the single-turn case, except that the user needs to define and register a multi-turn workflow class (by inheriting the base class MultiTurnWorkflow), where the logic of agent-environment interaction is implemented. Listing 2 provides one such example for ALFWorld [25]. For training efficiency, the process_messages_to_experience() method concatenates multiple rounds of agent-environment interactions compactly into an Experience instance consisting of a single token sequence with proper masking, which can readily be consumed by standard RL algorithms like PPO and GRPO.
@@ 4.2 Multi-Turn Workflow
18 response_text = self.get_model_response_text(memory)

### v2

@@ 2.3.1 End-to-end Data Transformation
  • The Formatter Module unifies disparate data sources into RFT-compatible formats, providing convenient conversion between raw inputs (e.g., meta-prompts, domain-specific corpora, and QA pairs with tagged rewards) and structured RFT representations. For efficient RFT workloads, we utilize buffer-based persistent storage (Section 2.1) to support different data models, such as ExperienceModel for prioritized rollout trajectories and DPODataModel for preference pairs. The conversion logic and data models are highly customizable to meet diverse requirements for managing experience data. This flexibility enables robust metadata recording and field normalization, which is essential for advanced scenarios such as asynchronous RFT in trainer-explorer environments, agent self-evolution from a cold start using meta-prompts, and knowledge injection from structurally complex domain-specific corpora.
  • The Controller Module manages the complete data pipeline lifecycle through distributed server initialization, declarative configuration, and automated dataset persistence. It implements dynamic control mechanisms for asynchronous scenarios and protection against resource exhaustion, with configurable termination conditions based on compute quota or data quantity. This modular design enables Trinity-RFT to handle data transformations flexibly while maintaining consistency across different RFT modes.
@@
Figure 5: The interaction of data processor and data buffers in Trinity-RFT, divided into two key stages. Left: Task Curation & Prioritization prepares the initial tasks for the explorer. Right: Experience Shaping processes the collected trajectories from the explorer before they are used by the trainer. The data processor is a central component that operates on different buffers at different stages.
2.3.2 Task Curation and Prioritization
Before the RFT loop begins, it is crucial to prepare a high-quality set of initial tasks. This stage, depicted on the left side of Figure 5, transforms raw data into an optimized task set for the explorer.
@@ 2.3.2 Task Curation and Prioritization
The process begins with raw data sources (e.g., prompts, domain corpora), which are ingested into a buffer. The Data Processor, powered by over 100 operators from Data-Juicer [2], reads from this buffer to perform various curation tasks. It provides composable building blocks for experience cleaning (e.g., length filters, duplication removal), safety alignment (e.g., toxicity detection, ethics checks), and preference data synthesis (e.g., critique-conditioned augmentation). By treating Data-Juicer as a modular data processing operator pool rather than a central dependency, Trinity-RFT provides RL-specific abstractions and coherence, while benefiting from well-established data tools.
The processed data is then organized into a structured task buffer. This stage effectively implements a form of curriculum learning by allowing users to prioritize tasks (e.g., from easy to hard), guiding the explorer towards a more efficient and stable learning trajectory from the outset. This entire workflow is managed by a service-oriented architecture that decouples data logic from procedural control, ensuring flexibility and scalability, especially in asynchronous and distributed settings.
2.3.3 Active Experience Shaping
Once the explorer begins interacting with the environment, it generates a continuous stream of experience data. To maximize learning efficiency, this raw experience must be actively shaped before it reaches the trainer. This stage is shown on the right side of Figure 5.
Generated experiences are first collected in a buffer. The Data Processor is applied again with a series of transformations to clean, augment, or synthesize these experiences. This is where the core of RFT data intelligence lies. Key capabilities include:
  • Agent-Driven Data Processing: Trinity-RFT introduces a powerful agentic paradigm for data manipulation. Users can define complex processing pipelines through high-level objectives, specified as either natural language commands (e.g., “improve safety” or “increase response diversity”) or explicit Data-Juicer configurations. The framework automatically translates these commands into executable workflows backed by its modular components like DataCleaner and DataSynthesizer. This design provides a user-friendly abstraction layer over the underlying Data-Juicer operators, making advanced processing functionalities accessible to both RFT users familiar with Data-Juicer and those who are not. It also facilitates the flexible injection of user-defined inductive biases into the learning process, unlocking new research directions for self-evolving agents, as we will discuss later in Section 2.3.5.
  • Online Reward Shaping: The data processor can dynamically augment the reward signal. Instead of relying on a single, often sparse, task-completion reward, users can add dense rewards based on quality, diversity, or safety scores computed on the fly. This enriched feedback provides a much stronger learning signal for the trainer.
  • Prioritized Experience Replay: Experiences are not treated equally. Trinity-RFT allows for flexible, multi-dimensional utility scoring to prioritize the most valuable samples for training. The DataActiveIterator supports version-controlled experience reuse and cross-task data lineage tracking, ensuring that the trainer always learns from the most informative data available. This mechanism is also critical for handling delayed rewards, as experience utilities can be updated asynchronously as new feedback arrives.
@@ 2.3.4 Human-AI Collaboration
The decoupled design of Trinity-RFT, and the presence of a standalone experience buffer in particular, enable human feedback to participate in RL loops without breaking the asynchronous execution model. For instance, human-verified samples can be prioritized for training while fresh experiences are being collected, which is a critical capability for real-world deployment scenarios with mixed feedback sources. Further details for human-AI collaboration in Trinity-RFT will be illustrated in Section 3.5.
2.3.5 Discussion: Unlocking New Research & Development Directions
The modular design of our data pipelines and the powerful data processor open up promising research and development avenues to be further explored.
One direction is about effective management of experience data. While prior RFT works often treat the experience as a static log, Trinity-RFT enables a more sophisticated, full-lifecycle approach to data, from selective acquisition to efficient representation:
  • Intelligent Perception and Collection: In an open-ended environment, what experience is “worth” recording? Storing everything creates a low signal-to-noise ratio and burdens the trainer. Trinity-RFT’s architecture allows researchers to implement active collection strategies. For instance, one could design a data processor operator that evaluates incoming experiences from the explorer based on metrics like surprise, uncertainty, or information gain, and only commits the most salient trajectories to the replay buffer. This transforms data collection from passive logging into a targeted, intelligent process.
  • Adaptive Representation: Raw experience is often high-dimensional and redundant (e.g., long dialogues, complex code generation traces). How can this be distilled into a format that an agent can efficiently learn from? The data processor in Trinity-RFT acts as a powerful transformation engine. Researchers can use it to explore various representation learning techniques, such as automatically summarizing trajectories, extracting causal chains from tool usage, or converting a multi-turn dialogue into a structured preference pair. This not only makes training more efficient but also opens the door to building meta-experience (more abstract and reusable knowledge) from raw interaction data.
  • Agentic Workflows: Trinity-RFT’s agent-driven processing enables the research and development of self-improving agents, e.g., by configuring the policy agent to also serve as the “processing agent” for LLM-based Data-Juicer operators. Such an agent could perform its own critique and dynamically curate its own training data, creating a truly autonomous learning and data management loop.
Another direction is about synthetic and counterfactual experience processing. The integration of synthesis operators enables research into creating “better-than-real” data. Instead of relying solely on the agent’s own trial-and-error, our framework facilitates exploring questions like:
  • Dynamic and Composable Rewarding: With our framework, researchers can move beyond static, hand-crafted rewards. It is now possible to investigate dynamic reward shaping, where auxiliary signals like novelty, complexity, or alignment scores are automatically extracted from trajectories and composed into a dense reward function. How to define “good” experience and how can we learn the optimal combination of these reward components as the agent’s policy evolves?
  • Experience Reorganization: Can successful sub-trajectories from different tasks be “spliced” together to solve a novel, composite task? For example, can an agent that has learned to “open a door” and “pick up a cup” synthesize a new trajectory to "enter the room and fetch the cup"?
  • Failure Repair: Can the data processor identify where errors occur in a failed trajectory, and synthesize a corrected version for the trainer to learn from, effectively turning failures into valuable lessons?
  • Success Amplification: Can a single successful experience be augmented into multiple diverse yet successful variants, thereby improving the generalization and robustness of the learned policy?
By providing dedicated capabilities for such advanced data and reward manipulation, Trinity-RFT aims to facilitates flexible processing of “experience data” for the next generation of self-evolving LLMs.
2.4 User-Friendliness
Trinity-RFT has been designed with user-friendliness as a top priority.
For development and research:
@@ 2.4 User-Friendliness
The modular and decoupled design of Trinity-RFT allows users to develop a new algorithm for a specific aspect of RFT by adding one or a few new classes that implement the essential functionalities of interest, without concerning about other aspects of RFT or intrusive modifications of the original codebase. In addition, we include a monitor (built upon Wandb [40] and TensorBoard [37]) that makes it easy to track the progress of an RFT process, both quantitatively (e.g., via learning curves for rewards and other metrics) and qualitatively (e.g., via concrete examples of rollout trajectories generated at different RL steps). See Figure 6 for an example snapshot of the monitor.
For RFT applications:
Trinity-RFT offers extensive graphical user interfaces to support low-code usage of the framework, and to maximize transparency of the RFT process. For example, we implement a configuration manager, as shown in Figure 7, that allows the user to create configuration files conveniently via a front-end interface. We also provide Trinity-Studio, an all-in-one unified UI (including the aforementioned monitor and configuration manager) that allows the user to configure and run data inspection, data processing, RFT learning process, etc., all by clicking the mouse and filling forms, without writing any code. An example for using Trinity-Studio will be introduced in Section 3.6. Such functionalities, of course, can be useful not only for applications but also for development and research.
@@
3 Examples, Applications, and Experiments
This section demonstrates the utilities and user-friendliness of Trinity-RFT and exemplifies some concepts introduced in previous sections, through a diverse range of examples, applications and experiments. Additional step-by-step tutorials can be found on the documentation website [^2] https://modelscope.github.io/Trinity-RFT, or the examples folder of the GitHub repository [^3] https://github.com/modelscope/Trinity-RFT/tree/main/examples.
3.1 Customizing Agent-Environment Interaction
With a modular design, Trinity-RFT can be easily adapted to a new downstream scenario by implementing the logic of agent-environment interaction in a single workflow class, without modifications to other components of the codebase. This approach is also sufficient for macroscopic RL algorithm design that targets high-quality experience synthesis with environmental feedback [4]. We provide some concrete examples in the rest of this subsection.
3.1.1 Single-turn Scenarios
In a simple yet common scenario, a user of Trinity-RFT would like to train an LLM for completing single-turn tasks, where the LLM generates one response to each input query. For this purpose, the user mainly needs to (1) define and register a single-turn workflow class (by inheriting the base class Workflow) tailored to the targeted tasks, and (2) specify the tasksets (for training and/or evaluation) and the initial LLM, both of which are compatible with HuggingFace [13] and ModelScope [18] formats.
Listing 1 gives a minimal example for implementing a single-turn workflow. Suppose that each task is specified by a <question, answer> tuple. The run() method of ExampleWorkflow calls the LLM once to generate a response for the question, calculates its reward, and returns an Experience instance that consists of the response itself, the reward value, and the log-probabilities of response tokens predicted by the rollout model (which is necessary for certain RL algorithms, such as PPO [27] and GRPO [28]). Some built-in workflows and reward functions have been implemented in Trinity-RFT, e.g., the MathWorkflow class for math-related tasks.
In some cases, the user wants to utilize auxiliary LLMs in the workflow, e.g., for computing rewards via LLM-as-a-judge, or for playing the roles of other agents in a multi-agent scenario. For these purposes, the user can specify auxiliary_models via APIs when initializing the workflow.
@@ 3.1.1 Single-turn Scenarios
4 def __init__(
5 self,
6 model: ModelWrapper,
7 task: Task,
8 auxiliary_models: Optional[List[openai.OpenAI]] = None,
9 ):
10 super().__init__(model, task, auxiliary_models)
@@
13
14 def calculate_reward_by_rule(self, response: str, truth: str) -> float:
@@
16
17 def calculate_reward_by_llm_judge(self, response: str, truth:str) -> float:
18 judge_model = self.auxiliary_models[0]
19 PROMPT_FOR_JUDGE = """Please evaluate..."""
20 completion = judge_model.chat.completions.create(
21 model="gpt-4", # Or another suitable judge model
22 messages=[{"role": "user", "content": PROMPT_FOR_JUDGE}],
23 )
24 reward_str = completion.choices[0].message.content.strip()
25 reward = float(reward_str)
26 return reward
27
@@
36 **self.rollout_args,
@@
38 reward: float = self.calculate_reward_by_rule(response.response_text, self.answer)
39 # reward: float = self.calculate_reward_by_llm_judge(response.response_text, self.answer)
@@ 41 Experience(
3.1.2 Multi-turn Scenarios
In more advanced cases, the user would like to train an LLM-powered agent that solves multi-turn tasks by repeatedly interacting with the environment. In Trinity-RFT, achieving this is mostly as simple as in the single-turn case, except that the user needs to define and register a multi-turn workflow class by inheriting the base class MultiTurnWorkflow. Listing 2 provides one such example using the ALFWorld dataset [30]. For training efficiency, the process_messages_to_experience() method concatenates multiple rounds of agent-environment interactions compactly into an Experience instance consisting of a single token sequence with proper masking, which can readily be consumed by standard RL algorithms like PPO and GRPO.
For more detailed examples of multi-turn cases, please refer to the documentation [^4] https://modelscope.github.io/Trinity-RFT/tutorial/example_multi_turn.html.
@@ 3.1.2 Multi-turn Scenarios
16 response_text = self.model.chat(memory, n=1)[0].response_text